### v13

# On $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ and Frege Systems

Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The main techniques used in this paper are lazy-diagonalization and the novel new technique developed in the author’s recent work [Lin21]. Further, since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mystery lies behind it and show that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is not enumerable, thus showing that the technique of lazy-diagonalization is not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there exists a language $L_{inter}\in{\rm co}\mathcal{NP}$, which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences implied by our main outcome, such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$, $\mathcal{BPP}\neq\mathcal{NEXP}$ and there exists no super proof system. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomially bounded.
1. Introduction
2. Preliminaries
3. All ${\rm co}\mathcal{NP}$-machines are enumerable
4. Lazy-diagonalization against All ${\rm co}\mathcal{NP}$-machines
5. Proving That $L_{d}\in\mathcal{NP}$
6. Breaking the Relativization Barrier
7. Rich Structure of ${\rm co}\mathcal{NP}$
8. Frege Systems
9. Conclusions
A Frege proof system $F$ is an inference system for propositional logic based on
  1. (1)
    a language of well-formed formulas obtained from a numerable set of propositional variables and any finite propositionally complete set of connectives;
  2. (2)
    a finite set of axiom schemes;
  3. (3)
    and the rule of Modus Ponens
    $\frac{A\quad A\rightarrow B}{B}.$
A proof $P$ of the formula $A$ in a Frege system is a sequence $A_{1},\cdots,A_{n}$ of formulas such that $A_{n}$ is $A$ and every $A_{i}$ is either an instance of an axiom scheme or it is obtained by the application of the Modus Ponens from premises $A_{j}$ and $A_{k}$ with $j,k<i$.
Remark 8.1
The well-known inference rule of Modus Ponens is its only rule of inference:
$\frac{A\quad A\rightarrow B}{B}$
We give a more intuitive explanation about the inference rule of Modus Ponens: For example, in the proof of Theorem 5.3 in Section 5 where we prove that $L_{d}\in\mathcal{NP}$, $A$ is the following:
$A\overset{\rm def}{=}(5.1)\bigwedge(5.2)\bigwedge(5.3),$
$B$ represents the proposition $L_{d}\in\mathcal{NP}$. Namely,
$B\overset{\rm def}{=}L_{d}\in\mathcal{NP},$
and one of the main deductions of Theorem 5.3, besides showing that $A$ is true, is to show that $A$ implies $B$.
The axiom schemes of a Frege proof system is the following (see e.g. [Bus99]):
$(P\wedge Q)\rightarrow P$
$(P\wedge Q)\rightarrow Q$
$P\rightarrow(P\vee Q)$
$Q\rightarrow(P\vee Q)$
$(P\rightarrow Q)\rightarrow((P\rightarrow\neg Q)\rightarrow\neg P)$
($(\neg\neg P)\rightarrow P$
$P\rightarrow(Q\rightarrow P\wedge Q)$
$(P\rightarrow R)\rightarrow((Q\rightarrow R)\rightarrow(P\vee Q\rightarrow R))$
$P\rightarrow(Q\rightarrow P)$
$(P\rightarrow Q)\rightarrow(P\rightarrow(Q\rightarrow R))\rightarrow(P\rightarrow R).$
More generally, a Frege system is specified by any finite complete set of Boolean connectives and finite set of axiom schemes and rule schemes, provided it is implicationally sound and implicationally complete.
Definition 8.3
The length of a Frege proof is the number of symbols in the proof. The length $|\varphi|$ of a formula $\varphi$ is the number of symbols in $\varphi$.
Now, we introduce the notion of $p$-simulation between two proof systems.
Definition 8.4 ([CR79])
If $f_{1}:\Sigma_{1}^{*}\rightarrow L$ and $f_{2}:\Sigma_{2}^{*}\rightarrow L$ are proof system for $L$, then $f_{2}$ $p$-simulates $f_{1}$ provided there is a function $g:\Sigma_{1}^{*}\rightarrow\Sigma_{2}^{*}$ such that $g\in\mathfrak{F}$, and $f_{2}(g(x))=f_{1}(x)$ for all $x$.
The notion of an abstract propositional proof system is given as follows:
Definition 8.5 ([Bus99])
An abstract propositional proof system is a polynomial-time computable function $g$ such that $Range(g)=TAUT$; i.e., the range of $g$ is the set of all Boolean tautologies. An $g$-proof of a formula $\phi$ is a string $w$ such that $g(w)=\phi$.
8.1 Lower Bounds for Frege Systems
Let $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ be two arbitrary Frege systems, the following theorem indicates that these standard proof systems for the propositional calculus are about equally powerful:
Theorem 8.1 ([Rec76, CR79])
Any two Frege systems $p$-simulate each other. Hence one Frege system is polynomially bounded if and only if all Frege systems are.
The following theorem gives a necessary and sufficient condition for the complexity class $\mathcal{NP}$ being closed under complementation. For completeness, the proof is also quoted from [CR79]:
Theorem 8.2 ([CR79])
$\mathcal{NP}$ is closed under complementation if and only if $TAUT$ is in $\mathcal{NP}$.
Proof. The “if” part. Suppose that the set of tautologies is in $\mathcal{NP}$. Then every set $L$ in $\mathcal{NP}$ is reducible to the complement of the tautologies [Coo71], i.e., there is a function $f$ that can be computed by a deterministic Turing machine in time bounded by a polynomial in the length of the input such that for all strings $x$, $x\in L$ iff $f(x)$ is not a tautology. Hence a nondeterministic procedure for accepting the complement of $L$ is: on input $x$, compute $f(x)$, and accept $x$ if $f(x)$ is a tautology, using the nondeterministic algorithm for tautologies assumed above. Hence the complement of $L$ is in $\mathcal{NP}$.
The “only if” part. Suppose that $\mathcal{NP}$ is closed under complementation. The the complement of the set of tautologies is in $\mathcal{NP}$, since to verify that a formula is not a tautology one can guess at a truth assignment and verify that it falsifies the formula.
Since nondeterministic Turing machines can simulate Frege proof (see e.g. [Bus99]) and by Theorem 8.1 we know that any two Frege systems $p$-simulate each others, we then can present the proof of Theorem 8 as follows:
Proof of Theorem 8. We show the theorem by contradiction. Suppose that there is a polynomial $p(n)$ such that for all $\psi\in TAUT$, there exists a proof $\oint$ of $\psi$ of length at most $p(|\psi|)$ (see Definition 8.3), then a nondeterministic Turing machine $M$ on input $\psi$ can guess $\oint$ of length at most $p(|\psi|)$ and checking whether $\oint$ is correct in deterministic polynomial time, and we know that the set of tautologies is ${\rm co}\mathcal{NP}$-complete, which, together with Theorem 8.2, implies
$\mathcal{NP}={\rm co}\mathcal{NP},$
a contradiction to Theorem 1. This finishes the proof.
Remark 8.2
Initially, propositional proof complexity has been primarily concerned with proving lower bounds (even conditional) for the length of proofs in propositional proof systems, which is extremely interesting and well justified in its own right, with the ultimate goal of settling whether $\mathcal{NP}={\rm co}\mathcal{NP}$ (see e.g. [CR79]). In fact, as mentioned in Section 1, such research directions are called Cook’s Program for separating $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (see e.g. [Bus12]): Prove superpolynomial lower bounds for proof lengths in stronger and stronger propositional proof systems until they are established for all abstract proof systems. However, our approach is “to do the opposite”, that is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems.
9 Conclusions
In conclusion, we have shown that
$\mathcal{NP}\neq{\rm co}\mathcal{NP},$
thus resolving a very important and long-standing conjecture in computational complexity theory. Specifically, we first showed that the set of all ${\rm co}\mathcal{NP}$-machines is enumerable. Then, we constructed a universal nondeterministic Turing machine $D$ to simulate all ${\rm co}\mathcal{NP}$-machines and to do the lazy-diagonalization operation to create a language $L_{d}$ which is not in ${\rm co}\mathcal{NP}$. Next, we applied the technique developed in the author’s recent work [Lin21] to show that the language $L_{d}$ accepted by the universal nondeterministic Turing machine $D$ is in fact in $\mathcal{NP}$, thus establishing the desired main conclusion of this paper.
We stress again that one of the important techniques used to prove lower bounds (i.e., Theorem 4.1) is the so-called lazy-diagonalization [FS07, AB09], and the essential technique used to show the upper bounds result, i.e., the result that the language $L_{d}\in\mathcal{NP}$, is essentially the same as the one developed in the author’s recent work [Lin21].
We have broken the so-called Relativization Barrier. Since by the result of [BGS75], it is not hard to show that there exists some oracle $A$ such that
$\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A},$
which suggests that (in the light of the point of view given in [BGS75]) ordinary diagonalization techniques (lazy-diagonalization) are not capable of proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$; see [BGS75]. We explored the mystery behind this phenomenon by proving that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}$-machines with oracle $A$ is in fact not enumerable, thus showing that the ordinary diagonalization techniques (i.e., lazy-diagonalization) will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
As a by-product, the important result that $\mathcal{P}\neq\mathcal{NP}$ which is shown in [Lin21], also follows. The above outcome can be seen as, in a sense, the direct or indirect consequence of the proof procedures of
$\mathcal{NP}\neq{\rm co}\mathcal{NP}.$
Note again that by Corollary 3 we in fact have shown $\mathcal{P}\neq{\rm co}\mathcal{NP}$ because if $\mathcal{P}={\rm co}\mathcal{NP}$, then $\mathcal{P}={\rm co}\mathcal{P}=\mathcal{NP}$, a contradiction to Corollary 3. As a matter of fact and interestingly, we have also shown by Theorem 6 that there exists a ${\rm co}\mathcal{NP}$-intermediate language in ${\rm co}\mathcal{NP}$, which is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete, i.e., the complexity class ${\rm co}\mathcal{NP}$ also has a rich structure. We have also obtained other interesting consequences, such as Corollary 4 (i.e., $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$), Corollary 5 (i.e., $\mathcal{BPP}\neq\mathcal{NEXP}$) and Corollary 7, which states that there is no super proof system.
As mentioned in the introduction section, the problem of $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ closely connects with the field of proof complexity. In particular, we have shown that no Frege proof systems are polynomially bounded (Theorem 8), thus answering an open problem in [Pud08]. As pointed out in Remark 8.2, we do the opposite direction of Cook’s program. That is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems. Note that one of the key points to show the Theorem 8 is that nondeterministic Turing machines can simulate Frege proofs (see [Bus99]). But we do not know whether the nondeterministic Turing machines can simulate Extended Frege proofs. For more details about Extended Frege Systems, please consult the references [CR79, CN10, Kra95].
Lastly, there are many other interesting open problems in the field of proof complexity that we did not touch on in this paper, and the interested reader is referred to these references [Bus12, CR74, CR79, CN10, Kra95, Kra19, Pud08, Raz03, Raz15] for those excellent open problems.